OpenStudy (anonymous):

lim x->0 ((1+x)/(2+x))*(1-x^1/2)/(1-x) can I just put 0 inside and get: 1/2 or there is something tricky here?

OpenStudy (anonymous):

U can Just Put 0 inside :)

OpenStudy (anonymous):

1/2*1/2 = 1/4 u should get

OpenStudy (anonymous):

I don't see anything preventing you from doing that, so yeah.

OpenStudy (anonymous):

@ganibl It should be 1/2, the second part should yield 1/1 instead of 1/2

OpenStudy (anonymous):

opops sorry my bad

OpenStudy (anonymous):

so, it's 1/2, thank you!

OpenStudy (anonymous):

now I see what's the point of the question: the next one is the same but x->1 And now it will be impossible to just insert 0.

OpenStudy (anonymous):

It's only necessary to do something "tricky", as you call it, if you get either something over 0 or an indeterminate form (like 0/0, 1^infinity, stuff like that)

OpenStudy (anonymous):

Well, actually, probably not even something over 0. I forgot his is already calculus.

OpenStudy (anonymous):

now I will need to use the principle with e?

OpenStudy (anonymous):

Easiest way I can think of is multiply them together first then apply l'hopital's rule. Not entirely sure if this will get you something workable, though

OpenStudy (anonymous):

I don't know l'hopital's rule, unfortunately.

OpenStudy (anonymous):

I mean, I am supposed to get the result using something else.

OpenStudy (anonymous):

Was that not part of your curriculum?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ah, I see. In that case, I need to look at it some more

OpenStudy (anonymous):

Just to be clear, is the question supposed to say 1-x^1/2 or (1-x)^1/2?

OpenStudy (anonymous):

1-x^1/2

OpenStudy (anonymous):

2+x/2+x + (-1)/2+x ?

OpenStudy (anonymous):

Multiply the second part by (1+x^1/2)/(1+x^1/2), you should be able to get rid of 1-x terms

OpenStudy (anonymous):

still got stuck...

OpenStudy (anonymous):

Working with the second part: (1-x^1/2)/(1/x) =((1-x^1/2)/(1/x))*((1+x^1/2)/(1+x^1/2)) =((1-x^1/2)(1+x^1/2))/((1-x)(1+x^1/2)) =(1-x)/((1-x)(1+x/^1/2)) =1/(1+x^1/2)

OpenStudy (anonymous):

After which there is nothing stopping you from substituting x=1 in.

OpenStudy (anonymous):

got it! thanks!